An astronaut has left an orbiting space shuttle to test a new personal maneuvering unit. As she moves along a straight line , her partner on board the shuttle measures her velocity every 2.0s starting at time t=1.0s as follows:


T

1.0s

3.0s

5.0s

V

0.8m/s

1.2m/s

1.8m/s


Find the average acceleration for each time interval:

a) 1s to 3s and b) 3s to 5s


1
Expert's answer
2021-03-29T09:03:55-0400

(a)

"a_{avg}=\dfrac{v_2-v_1}{t_2-t_1},""a_{avg}=\dfrac{1.2\ \dfrac{m}{s}-0.8\ \dfrac{m}{s}}{3.0\ s-1.0\ s}=0.2\ \dfrac{m}{s^2}."

(b)

"a_{avg}=\dfrac{1.8\ \dfrac{m}{s}-1.2\ \dfrac{m}{s}}{5.0\ s-3.0\ s}=0.3\ \dfrac{m}{s^2}."